### v2

@@ 4.2 Convergence Rates and Asymptotic Normality
Since I let $n\rightarrow\infty$ for each fixed $t$ in the asymptotic analysis, the $t$-subscript is suppressed for brevity of exposition when there is no confusion. Let $p_{i}\coloneqq p^{M_{2}}\left(X_{i},V_{i},W_{i}\right)$ and $P\coloneqq\mathbb{E}p_{i}p_{i}^{{}^{\prime}}$. As Imbens and Newey (2009) have studied the convergence rates for the series estimators of conditional CDF (e.g., $F_{\left.X\right|Z}$) and conditional expectation functions (e.g., $\mathbb{E}\left[\left.Y\right|X,V\right]$), I adapt to their analysis to prove convergence rates for $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$. I impose the following assumption.
@@
Let $\zeta\left(m_{2}\right)\coloneqq m_{2v}^{\theta}m_{2}$ and $\zeta_{1}\left(m_{2}\right)\coloneqq m_{2v}^{\theta+2}m_{2}$. With Assumption 5 in position, the next lemma follows directly from Theorem 12 of Imbens and Newey (2009).
@@
$\mathbb{E}\left[n^{-1}\sum_{i}\left(\widehat{V}_{i}-V_{i}\right)^{2}\right]$ $=O\left(n^{-1}M_{1}+M_{1}^{1-2d_{1}/j_{1}}\right)\eqqcolon O\left(\Delta_{1n}^{2}\right),$
[…]
$\sup_{s\in\mathcal{S}}\left|\widehat{G}\left(s\right)-G\left(s\right)\right|$ $=O_{p}\left(\zeta\left(m_{2}\right)\Delta_{2n}\right).$
Lemma 1 states that the mean-square convergence rate for $\widehat{G}\left(s\right)$ is the sum of the first-step rate $\Delta_{1n}^{2}$, the variance term $n^{-1}m_{2}$, and the squared bias term $m_{2}^{-2d_{2}/j_{2}}$. $d_{1}/j_{1}$ and $d_{2}/j_{2}$ are the uniform approximation rates that govern how well the unknown functions $V\left(x,z,w\right)$ and $G\left(s\right)$ can be approximated with $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$, respectively; see Assumption 3 and 5 of Imbens and Newey (2009) for the details. Note that even though the order of the basis function for the second-step estimation is $M_{2}$, by the index structure of (2.1) $M_{2}=d_{X}\times m_{2}$ and $d_{X}$ is finite. Thus, the effective order that matters for the convergence rate is $m_{2}$.
Let $\xi_{i}\coloneqq b_{1}\left(V_{i},W_{i}\right)-b\left(X_{i}\right)$ and $\xi\coloneqq\left(\xi_{1},...,\xi_{n}\right)^{{}^{\prime}}$. To derive the convergence rates of $\widehat{\overline{b}}$ and $\widehat{b}\left(x\right)$, I impose the following assumption.
@@
    There is a constant $C>0$ and $\zeta\left(M_{3}\right)$, such that for each $M_{3}$ there exists a non-singular constant matrix $N_{r}$ such that $\widetilde{r}^{M_{3}}\left(x\right)\coloneqq N_{r}r^{M_{3}}\left(x\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{r}^{M_{3}}\left(X_{i}\right)\widetilde{r}^{M_{3}}\left(X_{i}\right)^{{}^{\prime}}\right)\geq C$ and $\sup_{x\in\mathcal{X}}\left\|\widetilde{r}^{M_{3}}\left(x\right)\right\|\leq C\zeta\left(M_{3}\right)$.
@@
    $\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]$ is bounded.
Assumption 6 imposes conditions on the degree of smoothness of $b\left(x\right)$, the normalization of basis functions $r^{M_{3}}\left(x\right)$, and the boundedness of the second moment of $\xi_{i}$, similar to those in Assumption 5. Since $\widehat{b}_{1}\left(v,w\right)$ and $\widehat{G}\left(s\right)$ share the same series regression coefficient $\widehat{\alpha}^{M_{2}}$, the convergence rate of $\widehat{b}_{1}\left(v,w\right)$ to $b_{1}\left(v,w\right)$ is the same as that of $\widehat{G}\left(s\right)$ to $G\left(s\right)$. I use this result to prove the convergence rates of $\widehat{\overline{b}}$ and $\widehat{b}\left(x\right)$, both of which are unknown but estimable functionals of $\widehat{G}\left(s\right)$.
@@
[…]
$\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|$ $=O_{p}\left(\zeta\left(M_{3}\right)\Delta_{3n}\right).$
The 2002 working paper version of Imbens and Newey (2009) (henceforth IN02) has obtained asymptotic normality for estimators of known and scalar-valued linear functionals of $G\left(s\right)$. I adapt to their results to analyze vector-valued functionals of $G\left(s\right)$ via a Cramér–Wold device and prove asymptotic normality for $\widehat{b}_{1}\left(v,w\right)$. With slight abuse of notation, I take $\widetilde{\overline{p}}^{M_{2}}\left(s\right)$ and $\widetilde{r}^{M_{3}}\left(x\right)$ that satisfy Assumptions 5 and 6 as $\overline{p}^{M_{2}}\left(s\right)$ and $r^{M_{3}}\left(x\right)$ in what follows.
@@
    There is a constant $C>0$ and $\zeta\left(M_{1}\right)$, such that for each $M_{1}$ there exists a non-singular constant matrix $N_{q}$ such that $\widetilde{q}^{M_{1}}\left(z,w\right)\coloneqq N_{q}q^{M_{1}}\left(z,w\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)^{{}^{\prime}}\right)\geq C$ and $\sup_{\left(z,w\right)\in\mathcal{Z\times\mathcal{W}}}\left\|\widetilde{q}^{M_{1}}\left(z,w\right)\right\|\leq C\zeta\left(M_{1}\right)$.
@@
    $G\left(s\right)$ is twice continuously differentiable with bounded first and second derivatives. For functional $a\left(\cdot\right)$ of $G$ and some constant $C>0$, it is true that $\left|a\left(G\right)\right|\leq C\sup_{s}\left|G\left(s\right)\right|$ and either (i) there is $\delta\left(s\right)$ and $\widetilde{\alpha}^{M_{2}}$ such that $\mathbb{E}\delta\left(S_{i}\right)^{2}<\infty$, $a\left(p_{m}^{M_{2}}\right)=\mathbb{E}\delta\left(S_{i}\right)p_{m}^{M_{2}}\left(S_{i}\right)$ for all $m=1,...,M_{2}$, $a\left(G\right)=\mathbb{E}\delta\left(S_{i}\right)G\left(S_{i}\right)$, and $\mathbb{E}\left(\delta\left(S_{i}\right)-p^{M_{2}}\left(S_{i}\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right)^{2}\rightarrow 0$; or (ii) for some $\widetilde{\alpha}^{M_{2}}$, $\mathbb{E}\left[p^{M_{2}}\left(S_{i}\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right]^{2}\rightarrow 0$ and $a\left(p^{M_{2}}\left(\cdot\right)^{{}^{\prime}}\widetilde{\alpha}^{M_{2}}\right)$ is bounded away from zero as $M_{2}\rightarrow\infty$.
@@
    $\sup\left\{\sup_{s\in\mathcal{S}}\left|G\left(s\right)-p^{M_{2}}\left(s\right)^{{}^{\prime}}\overline{\alpha}^{M_{2}}\right|,\sup_{s\in\mathcal{S}}\left|\partial\left(G\left(s\right)-p^{M_{2}}\left(s\right)^{{}^{\prime}}\overline{\alpha}^{M_{2}}\right)/\partial s_{j}\right|\right\}=O\left(M_{2}^{-d_{4}/j_{4}}\right).$
@@
[…]
Assumptions 7(a)(e) are also imposed by IN02 and are regularity conditions required for the asymptotic normality of $\widehat{b}_{1}\left(v,w\right)$. Assumption 7(f) concerns the asymptotic covariance matrix of $\widehat{b}_{1}\left(v,w\right)$ and is used by Andrews (1991). It guarantees that the normality result of IN02 applies to vector-valued functionals of $G\left(s\right)$. Essentially, Assumption 7(f) requires that all the coordinates of $\widehat{b}_{1}\left(v,w\right)$ converge at the same speed, which is a mild assumption in my setting because ex-ante I do not distinguish any coordinate of $\beta_{it}$ from the others.
@@
If the conditions of Theorem 2 and Assumption 7 are satisfied, then,
[…]
@@
Lemma 2 concerns $b_{1}\left(v,w\right)$, a known functional of $G\left(s\right)$. Therefore, the results of IN02 directly apply and I omit its proof in this paper. However, the results of IN02 do not directly apply to $\overline{b}$ and $b\left(x\right)$ because they are unknown functionals of $G\left(s\right)$. To explain, notice that by the LIE
$\overline{b}=\mathbb{E}\left[\partial G\left(S\right)/\partial X\right]\text{ and }b\left(x\right)=\mathbb{E}\left[\left.\partial G\left(S\right)/\partial X\right|X=x\right],$ (4.6)
@@
    $\mathbb{E}\overline{p}_{i}r_{i}^{{}^{\prime}}$ has full column rank.
@@
    $\mathbb{E}\left[\left.\left\|\xi\right\|^{4}\right|{\bf X}\right]$ is bounded and $\mathbb{E}\left[\left.\xi\xi^{{}^{\prime}}\right|{\bf X}\right]$ is bounded away from zero.
@@
[…]
Assumption 8(a) is needed to show that the asymptotic covariance matrix $\Omega_{2}$ of $n^{1/2}\left(\widehat{b}\left(x\right)-b\left(x\right)\right)$ is positive definite. Assumption 8(b) is a regularity condition imposed for the Lindeberg–Feller central limit theorem (CLT). Assumption 8(c) is similar to Assumption 7(f) and is needed to prove that the asymptotic normality result holds for vector-valued functionals of $G\left(s\right)$.
@@
where $\widehat{\Omega}_{2}$ is defined in (A.45).
Furthermore, if $\mathbb{E}\left\|b_{1}\left(v,w\right)-\overline{b}\right\|^{4}<\infty$, then
@@
where $\widehat{\Omega}_{3}$ is defined in (A.58).
5 Empirical Illustration
In this section, I apply my procedure to estimate a heterogeneous Cobb-Douglas production function for each of the five largest manufacturing sectors in China. I obtain unconditional means of the output elasticities and compare them with those derived using classic methods on the same data set. Furthermore, I estimate conditional means of the elasticities given the regressors. Results show that there are significant across-firm variations in the output elasticities. In Appendix C, I conduct a simulation study motivated by production function applications to support the findings of this illustration.

### v3

@@ 4.2 Convergence Rates and Asymptotic Normality
Since I let $n\rightarrow\infty$ for each fixed $t$ in the asymptotic analysis, the $t$-subscript is suppressed when it is clear. Let $p_{i}\coloneqq p^{M_{2}}\left(X_{i},V_{i},W_{i}\right)$ and $P\coloneqq\mathbb{E}p_{i}p_{i}^{\prime}$. In the analysis, I use results from Imbens and Newey (2009) to establish convergence rates for $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$. The following assumption is needed.
@@
Let $\zeta\left(m_{2}\right)\coloneqq m_{2v}^{\theta}m_{2}$ and $\zeta_{1}\left(m_{2}\right)\coloneqq m_{2v}^{\theta+2}m_{2}$. With Assumption 5 in place, the next lemma follows directly from Theorem 12 of Imbens and Newey (2009). Let $\Delta_{1n}^{2}:=n^{-1}M_{1}+M_{1}^{1-2d_{1}/j_{1}}$ and $\Delta_{2n}^{2}:=\Delta_{1n}^{2}+n^{-1}m_{2}+m_{2}^{-2d_{2}/j_{2}}$ in the next lemma.
@@
$\mathbb{E}\left[n^{-1}\sum_{i}\left(\widehat{V}_{i}-V_{i}\right)^{2}\right]$ $=O\left(\Delta_{1n}^{2}\right),$
$\int\left[\widehat{G}\left(s\right)-G\left(s\right)\right]^{2}dF\left(s\right)$ $=O_{p}\left(\Delta_{2n}^{2}\right),\quad\text{and}\quad\sup_{s\in\mathcal{S}}\left|\widehat{G}\left(s\right)-G\left(s\right)\right|=O_{p}\left(\zeta\left(m_{2}\right)\Delta_{2n}\right).$
Lemma 1 states that the mean-square convergence rate for $\widehat{G}\left(s\right)$ is the sum of the first-step rate $\Delta_{1n}^{2}$, the variance term $n^{-1}m_{2}$, and the squared bias term $m_{2}^{-2d_{2}/j_{2}}$. $d_{1}/j_{1}$ and $d_{2}/j_{2}$ are the uniform approximation rates that govern how well the unknown functions $V\left(x,z,w\right)$ and $G\left(s\right)$ can be approximated with $\widehat{V}\left(x,z,w\right)$ and $\widehat{G}\left(s\right)$, respectively; see Assumptions 3 and 5 of Imbens and Newey (2009) for the details.
Let $\xi_{i}\coloneqq b_{1}\left(V_{i},W_{i}\right)-b\left(X_{i}\right)$ and $\xi\coloneqq\left(\xi_{1},\ldots,\xi_{n}\right)^{\prime}$. To derive the convergence rates for the APE and LAR estimators, I impose the following assumption.
@@
    There is a constant $C>0$ and $\zeta\left(M_{3}\right)$, such that for each $M_{3}$ there exists a non-singular constant matrix $N_{r}$ such that $\widetilde{r}^{M_{3}}\left(x\right)\coloneqq N_{r}r^{M_{3}}\left(x\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{r}^{M_{3}}\left(X_{i}\right)\widetilde{r}^{M_{3}}\left(X_{i}\right)^{\prime}\right)\geq C$ and $\sup_{x\in\mathcal{X}}\left\|\widetilde{r}^{M_{3}}\left(x\right)\right\|\leq C\zeta\left(M_{3}\right)$.
@@
    $\mathbb{E}\left[\left.\xi\xi^{\prime}\right|{\bf X}\right]$ is bounded.
Assumption 6 imposes conditions on the degree of smoothness of $b\left(x\right)$, the normalization of basis functions $r^{M_{3}}\left(x\right)$, and the boundedness of the second moment of $\xi_{i}$, similar to those in Assumption 5. Since $\widehat{b}_{1}\left(v,w\right)$ and $\widehat{G}\left(s\right)$ share the same series regression coefficient $\widehat{\alpha}^{M_{2}}$, the convergence rate of $\widehat{b}_{1}\left(v,w\right)$ to $b_{1}\left(v,w\right)$ is the same as that of $\widehat{G}\left(s\right)$ to $G\left(s\right)$. I use this result to prove the convergence rates of $\widehat{\overline{b}}$ and $\widehat{b}\left(x\right)$, both of which are unknown but estimable functionals of $\widehat{G}\left(s\right)$. Let $\Delta_{3n}^{2}:=n^{-1}M_{3}+M_{3}^{-2d_{3}/j_{3}}+\Delta_{2n}^{2}$ in the next theorem.
@@
$\int\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|^{2}dF(x)$ $=O_{p}\left(\Delta_{3n}^{2}\right),\quad\text{and}\quad\sup_{x\in\mathcal{X}}\left\|\widehat{b}\left(x\right)-b\left(x\right)\right\|=O_{p}\left(\zeta\left(M_{3}\right)\Delta_{3n}\right).$
The 2002 working paper version of Imbens and Newey (2009) (henceforth IN02) has obtained asymptotic normality for estimators of known and scalar-valued linear functionals of $G\left(s\right)$. I build on their results to analyze vector-valued functionals of $G\left(s\right)$ via a Cramér–Wold device and prove asymptotic normality for $\widehat{b}_{1}\left(v,w\right)$. To simplify the notation, I take $\widetilde{\overline{p}}^{M_{2}}\left(s\right)$ and $\widetilde{r}^{M_{3}}\left(x\right)$ that satisfy Assumptions 5 and 6 as $\overline{p}^{M_{2}}\left(s\right)$ and $r^{M_{3}}\left(x\right)$ in what follows.
@@
    There is a constant $C>0$ and $\zeta\left(M_{1}\right)$, such that for each $M_{1}$ there exists a non-singular constant matrix $N_{q}$ such that $\widetilde{q}^{M_{1}}\left(z,w\right)\coloneqq N_{q}q^{M_{1}}\left(z,w\right)$ satisfies $\lambda_{\min}\left(\mathbb{E}\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)\widetilde{q}^{M_{1}}\left(Z_{i},W_{i}\right)^{\prime}\right)\geq C$ and $\sup_{\left(z,w\right)\in\mathcal{Z\times\mathcal{W}}}\left\|\widetilde{q}^{M_{1}}\left(z,w\right)\right\|\leq C\zeta\left(M_{1}\right)$.
@@
    $G\left(s\right)$ is twice continuously differentiable with bounded first and second derivatives. For functional $a$ of $G$ and some constant $C>0$, it is true that $\left|a\left(G\right)\right|\leq C\sup_{s}\left|G\left(s\right)\right|$ and either (i) there is $\delta\left(s\right)$ and $\widetilde{\alpha}^{M_{2}}$ such that $\mathbb{E}\delta\left(S_{i}\right)^{2}<\infty$, $a\left(p_{m}^{M_{2}}\right)=\mathbb{E}\delta\left(S_{i}\right)p_{m}^{M_{2}}\left(S_{i}\right)$ for all $m=1,...,M_{2}$, $a\left(G\right)=\mathbb{E}\delta\left(S_{i}\right)G\left(S_{i}\right)$, and $\mathbb{E}\left(\delta\left(S_{i}\right)-p^{M_{2}}\left(S_{i}\right)^{\prime}\widetilde{\alpha}^{M_{2}}\right)^{2}\rightarrow 0$; or (ii) for some $\widetilde{\alpha}^{M_{2}}$, $\mathbb{E}\left[p^{M_{2}}\left(S_{i}\right)^{\prime}\widetilde{\alpha}^{M_{2}}\right]^{2}\rightarrow 0$ and $a\left(p^{M_{2}}{}^{\prime}\widetilde{\alpha}^{M_{2}}\right)$ is bounded away from zero as $M_{2}\rightarrow\infty$.
@@
    $\sup\left\{\sup_{s\in\mathcal{S}}\left|G\left(s\right)-p^{M_{2}}\left(s\right)^{\prime}\overline{\alpha}^{M_{2}}\right|,\sup_{s\in\mathcal{S}}\left|\partial\left(G\left(s\right)-p^{M_{2}}\left(s\right)^{\prime}\overline{\alpha}^{M_{2}}\right)/\partial s_{j}\right|\right\}=O\left(M_{2}^{-d_{4}/j_{4}}\right).$
@@
    (Assumption J(iii) of Andrews (1991)) For a bounded sequence of constants $\left\{c_{1n}:n\geq 1\right\}$ and constant pd matrix $\overline{\Omega}_{1}$, it is true that $c_{1n}\Omega_{1}\overset{p}{\longrightarrow}\overline{\Omega}_{1}$, where $\Omega_{1}$ is defined in (A.12).
Assumption 7(a)(e) is also imposed by IN02 and is a regularity condition required for the asymptotic normality of $\widehat{b}_{1}\left(v,w\right)$. Assumption 7(f) concerns the asymptotic covariance matrix of $\widehat{b}_{1}\left(v,w\right)$ and is used by Andrews (1991). It guarantees that the normality result of IN02 applies to vector-valued functionals of $G\left(s\right)$. Essentially, Assumption 7(f) requires that all the coordinates of $\widehat{b}_{1}\left(v,w\right)$ converge at the same speed, which is mild because ex-ante I do not distinguish any coordinate of $\beta_{it}$ from the others.
@@
If the conditions of Theorem 2 and Assumption 7 are satisfied, then
$\sqrt{n}\,\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)\overset{d}{\longrightarrow}N\left(0,I\right),$
where $\Omega_{1}$ is defined in (A.12).
Furthermore,
$\widehat{\Omega}_{1}\Omega_{1}^{-1}\overset{p}{\longrightarrow}I,$
@@
Lemma 2 concerns $b_{1}\left(v,w\right)$, a known functional of $G\left(s\right)$. Therefore, the results of IN02 directly apply and I omit its proof in this paper. However, the results of IN02 do not directly apply to $\widehat{\overline{b}}$ and $\widehat{b}(x)$ because $\overline{b}$ and $b(x)$ are unknown functionals of $G\left(s\right)$. To explain, notice that by the law of iterated expectations
$\overline{b}=T^{-1}\sum_{t=1}^{T}\mathbb{E}\left[\partial G\left(S\right)/\partial X\right],\quad\text{and}\quad b\left(x\right)=\mathbb{E}\left[\left.\partial G\left(S\right)/\partial X\right|X=x\right],$ (4.8)
@@
    $\mathbb{E}\overline{p}_{i}r_{i}^{\prime}$ has full column rank.
@@
    $\mathbb{E}\left[\left.\left\|\xi\right\|^{4}\right|{\bf X}\right]$ is bounded and $\mathbb{E}\left[\left.\xi\xi^{\prime}\right|{\bf X}\right]$ is bounded away from zero.
@@
    For a sequence of bounded constants $\left\{c_{2n},c_{3n}:n\geq 1\right\}$ and some constant pd matrix $\overline{\Omega}_{2}$ and $\overline{\Omega}_{3}$, $c_{2n}\Omega_{2}\overset{p}{\longrightarrow}\overline{\Omega}_{2}$ and $c_{3n}\Omega_{3}\overset{p}{\longrightarrow}\overline{\Omega}_{3}$, where $\Omega_{2}$ and $\Omega_{3}$ are defined in (A.18) and (A.64), respectively.
  4. (d)
    $\mathbb{E}\left\|b_{1}\left(v,w\right)-\overline{b}\right\|^{4}<\infty$.
Assumption 8(a) is needed to show that the asymptotic covariance matrix $\Omega_{2}$ of $\sqrt{n}(\widehat{b}\left(x\right)-b\left(x\right))$ is positive definite and similarly for $\Omega_{3}$. Assumption 8(b) is a regularity condition imposed for the Lindeberg–Feller central limit theorem (CLT). Assumption 8(c) is similar to Assumption 7(f) and is needed to prove that the asymptotic normality result holds for vector-valued functionals of $G\left(s\right)$ by the Cramér–Wold device.
@@
$\sqrt{n}\,\Omega_{2}^{-1/2}(\widehat{b}\left(x\right)-b\left(x\right))$ $\overset{d}{\longrightarrow}N\left(0,I\right)\quad\text{and}\quad\sqrt{n}\,\Omega_{3}^{-1/2}(\widehat{\overline{b}}-\overline{b})\overset{d}{\longrightarrow}N(0,I),$
where $\Omega_{2}$ and $\Omega_{3}$ are defined in (A.18) and (A.64), respectively.
Furthermore,
$\widehat{\Omega}_{2}\Omega_{2}^{-1}\overset{p}{\longrightarrow}I\quad\text{and}\quad\widehat{\Omega}_{3}\Omega_{3}^{-1}\overset{p}{\longrightarrow}I,$
where $\widehat{\Omega}_{2}$ and $\widehat{\Omega}_{3}$ are defined in (A.48) and (A.67), respectively.
I present the main idea to the proof of Theorem 3, focusing on how the error propagates through the three-step analysis and the roles of $\widehat{\Omega}_{2}$ and $\widehat{\Omega}_{3}$. Define the functionals
$a\left(b_{1},V\right)$ $\coloneqq\mathbb{E}\left[\left.b_{1}\left(V,W\right)\right|X=x\right]=b\left(x\right),$
$\widehat{a}\left(b_{1},V\right)$ $:=\widehat{\mathbb{E}}\left[\left.b_{1}\left(V,W\right)\right|X=x\right],$
@@
Then
[…]
This decomposition makes clear that the asymptotic covariance of $\sqrt{n}\,(\widehat{b}(x)-b(x))$, denoted by $\Omega_{2}$ as in (A.18), must account for errors of all three sources. In the proof of Theorem 3, I show that
$\sqrt{n}\,\Omega_{2}^{-1/2}\big(\widehat{b}(x)-b(x)\big)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\sum_{k=1}^{3}\psi_{ki}+o_{p}(1),$ (4.10)
where the influence functions $\psi_{1i},\ \psi_{2i},\text{ and }\psi_{3i}$ correspond exactly to the three error terms on the right-hand side of (4.9) up to proper normalization. Once (4.10) is established, I verify the Lindeberg-Feller condition on $\Psi_{in}:=n^{-1/2}\sum_{k=1}^{3}\psi_{ki}$, which proves its convergence to the standard normal distribution and that $\Omega_{2}$ is indeed the appropriate asymptotic covariance associated with $\sqrt{n}\,(\widehat{b}(x)-b(x))$.
For feasible inference, a consistent estimator for $\Omega_{2}$ is needed, and I show in the proof of Theorem 3 that $\widehat{\Omega}_{2}$ is consistent. Note that $\Omega_{2}$ may diverge because $\widehat{b}(x)$ converges more slowly than $n^{-1/2}$ as shown in Theorem 2. Hence, all convergence results are expressed in self-normalized way, for example,
$\sqrt{n}\,\Omega_{2}^{-1/2}\big(\widehat{b}(x)-b(x)\big)\overset{d}{\longrightarrow}N(0,I)\quad\text{and}\quad\widehat{\Omega}_{2}\Omega_{2}^{-1}\overset{p}{\longrightarrow}I.$ (4.11)
The same argument applies to $\widehat{\Omega}_{3}$ with the basis function $r^{M_{3}}\left(x\right)=1$ and the influence function capturing serial correlation of estimating the APE across periods.
5 Empirical Application
In this section, I apply my procedure to estimate a heterogeneous Cobb-Douglas production function for each of the five largest manufacturing sectors in China. I obtain unconditional means of the output elasticities (APE) and compare them with those derived using classic methods on the same data set. Furthermore, I estimate conditional means of the elasticities given the regressors (LAR). Results show that there are significant across-firm variations in the output elasticities. I also examine how heterogeneity in output elasticities relates to observed characteristics and obtain intuitive results.
In Appendix B, I conduct several robustness checks—including alternative price deflators, sample trimming, alternative IV and $W_{i}$ constructions, city–year fixed effects, and comparison with OLS and naive IV estimates—and the results remain stable. In Appendix C, I conduct a production-function motivated simulation study to support the findings in this application.